# DySurv: Dynamic Deep Learning Model for Survival Analysis with Conditional Variational Inference

### v2

@@ Abstract
Survival analysis focuses on estimating time-to-event distributions which can help in dynamic risk prediction in healthcare. Extending beyond the classical Cox model, deep learning techniques have been developed which moved away from the constraining assumptions of proportional hazards. Traditional statistical models often only include static information where, in this work, we propose a novel conditional variational autoencoder-based method called DySurv, which uses a combination of static and time-series measurements from patient electronic health records to estimate the risk of death dynamically. DySurv has been tested on several time-to-event benchmarks where it outperforms existing methods, including deep learning methods, and we evaluate it on real-world intensive care unit data from MIMIC-IV and eICU. The predictive capacity of DySurv is consistent and the survival estimates remain disentangled across different datasets supporting the idea that dynamic deep learning models based on conditional variational inference in multi-task cases can be robust models for survival analysis.
@@ I Introduction
Survival analysis refers to statistical approaches to estimating distributions of event times or times it takes for an event to happen as well as rates of survival over time while accounting for censoring. The events in question can be machine failures in industry or the occurrence of specific diseases and death [1]. In clinical practice, survival analysis can play a key role and provide valuable information to predict patient outcomes and guide treatment decisions [2]. While most traditional applications of survival analysis occur in epidemiology at the population level, with the rise of deep learning techniques, personalised estimation of survival times for individual patients has become possible [3]. However, the limitations of standard statistical models such as the Cox proportional hazards model include lack of complexity, constraining assumptions about the behaviour or proportions of covariate effects over time, and the reliance on only using static covariates [4]. While there are extensions of Cox to time-varying covariates, these still suffer from the proportionality assumption (often violated in practice) and are often outperformed by simple deep learning models [5]. In settings like healthcare and with the occurrence of electronic health records, high-dimensional longitudinal information on the patient’s state can be an informative source for the prediction of clinical risk of mortality and other events.
@@
To incorporate longitudinal, high-dimensional, and potentially multimodal data and to move away from constraining assumptions of simple statistical models, deep learning methods have been proposed in survival analysis. These include extensions to the Cox model which estimate the parameters using a deep learning model that thus remain restrained with the proportionality assumption as in the case of DeepSurv [6]. Other deep learning models include DeepHit and its extension to longitudinal data, Dynamic-DeepHit, which implements a custom loss and a simple recurrent neural network to avoid making any parametric assumptions on the survival distribution [7, 8]. Simple extensions of recurrent neural networks for estimating survival distributions have also been proposed [9, 10]. More recently, the autoencoder structure has been proposed to learn from a combination of data modalities by finding a lower-dimensional latent representation of the data. ConcatAE uses a simple autoencoder on multi-omics data to find a hidden representation then used in a multi-class classification task to predict discrete survival times [11]. This treats the survival analysis task as one of simple classification and does not aim to estimate the survival distribution at all. On the other hand, the variational autoencoder present in VAECox, directly estimates the survival distribution using the Cox loss function while taking advantage of the variational inference over traditional autoencoders [12]. Due to the reliance on the Cox loss, VAECox suffers from the often-violated proportionality assumption meaning its results in real-world datasets could be limited.
@@
Our work addresses these limitations by using a cumulative incidence risk estimation loss function based on the negative log-likelihood which requires making no parametric or proportionality assumptions on the survival distribution and hazard risks. We also extend the variational setting to conditional VAE which we find improves predictive performance. Our model, DySurv, is compatible with both static and longitudinal time-series data enabling more comprehensive learning from patient electronic health records. We validate our approach both in static and time-varying settings using six benchmark datasets as well as two large public open-access ICU datasets. The critical care setting provides high-frequency longitudinal measurements and short-term health outcomes in which dynamic risk stratification of survival analysis could provide additional insights [13]. We aim to show that conditional variational inference and autoencoder reconstruction tasks can improve learning from complex time-series data by extracting latent features to optimise the survival task. Since DySurv is not only multimodal in input but also in output with both reconstruction and survival tasks included, conditional variational autoencoders (CVAE) provide better predictive performance [14]. By mapping high-dimensional covariates into a lower-dimensional space, CVAEs can effectively capture the underlying structure and identify hidden factors that significantly impact survival outcomes to predict mortality as well as time-to-death. This capability allows deeper insights into dynamic mortality risk prediction in healthcare.
II Materials and Methods
@@ II-A Data
Standard benchmark datasets contain only static features but here we implement survival analysis on the ICU datasets from MIMIC-IV and eICU which contain both static and time-series data. To show performance across different datasets and with different sizes, we will succinctly introduce these datasets. Across all datasets, the event in question is death. The datasets were split into 60% training, and 20% each for validation and testing. Quantile transformations have been applied for standardisation and fit only on the training dataset. Please see the appendix for a full list of features for each dataset.
@@ II-A1 SUPPORT
The Study to Understand Prognoses and Preferences for Outcomes and Risks of Treatments contains data from five care academic centres in the United States for patient survival in the following six months [15]. The result of the study was a prognostic model to estimate survival for seriously ill hospitalised patients. The dataset consists of 8,873 samples and 14 static features.
@@ II-A2 METABRIC
The Molecular Taxonomy of Breast Cancer International Consortium contains genetic and clinical data from breast cancer patients with 1,904 samples and 9 static features [16].
@@ II-A3 GBSG
The Rotterdam & German Breast Cancer Study Group contains treatment and clinical data on 2,232 breast cancer patients with 6 static features [17].
@@ II-A4 NWTCO
The National Wilm’s Tumor dataset contains staging and clinical data on 4,028 Wilms’ tumour patients with 6 static features [18].
@@
The simulated dataset contains discrete event times with 44 static features and 100,000 samples [19].
@@
The simulated dataset contains discrete event times with 5 static features and 50,000 samples [20].
@@ II-A7 MIMIC IV and eICU
We conduct experiments on the de-identified real-world ICU dataset Medical Information Mart for Intensive Care (MIMIC-IV v. 2.0, July 2022) including discharge information for more than 15,000 additional ICU patients compared to the previous release [21]. The dataset contains data from Beth Israel Deaconess Medical Center collected between 2008 and 2019. The dataset contains 71,935 samples of ICU stays with 33 static features (categorical features were one-hot encoded) and 65 time-varying features. Furthermore, we tested survival analysis models in the eICU Collaboration Research Database [pollard2018eicu]. The eICU database was processed using postgreSQL and the pandas package. eICU is a multi-centre ICU database with over 200,859 patient unit encounters for 139,367 unique patients admitted between 2014 and 2015 to one of 335 ICUs at 208 hospitals located throughout the United States. The database is de-identified and includes vital sign measurements, demographic data, and diagnosis information. Static variables include age, sex, admission unit, and others that did not have missingness. For the time-series variables, we use forward filling as clinicians in practice would only consider the last recorded measurement. If the first set of measurements is missing for some time-varying features, instead of dropping those features or patients, we backward fill from the closest measurement in the future. The processing of MIMIC-IV follows from previously published work on this dataset but is adapted for the survival scenario with the duration of stay in the ICU or the maximum time horizon for the event times defined as 10 days and time-series features taken in 72-hour timesteps [22]. Each node of the neural network model then corresponds to a 1-day risk prediction in the ICU as the maximum time horizon is 10 days. Censoring, as always, is defined as loss-to-follow-up or discharge.
@@ II-B DySurv
$f(t)=\lim_{\Delta t\rightarrow 0}\frac{\operatorname{P}(t<T\leq t+\Delta t)}{\Delta t}$
and $F(t)$ corresponds to the cumulative risk or incidence function $F(t)=\operatorname{P}(T\leq t)$. $T$ represents the time of the event, $\operatorname{P}$ is the probability, and $t$ is the specitic timestamp for risk estimation. The key to training a deep learning model is to learn an estimate of the cumulative incidence or risk function $\hat{F}(t)$ as the joint distribution of the event time and outcome label given the observations. In general, the time-to-event values can be left to be continuous depending on the model being considered but we discretise the time set into 10 equally spaced time periods for each dataset in the fashion of DeepHit and Dynamic-DeepHit [8]. As we discretise the time into intervals, we can estimate this event probability across arbitrary periods and remain faithful to the original survival analysis formulation rather than resorting to chained binary classification [23]. Furthermore, measurements are often right-censored, meaning that patients can leave the study or be lost to follow-up and they therefore do not experience an event at all. A common assumption is that this censoring is not important and independent of the outcome of the study itself [24]. Once we have an estimate of the risk, we can then simply obtain the survival function and curves by taking its negation from one as shown in equation (3).
@@
with $\mathcal{X}^{i}$ representing the feature matrix of patient $i$, $\mathcal{t}$ being the time-to-event ie. the minimum of the event and censoring time, $y$ being the label for the outcome, and $N$ samples included. A feature matrix (in the time-series case) for patient $i$, $\mathcal{X}^{i}$, can be seen as
$\mathcal{X}^{i}=\left\{\boldsymbol{x}_{1}^{i},\boldsymbol{x}_{2}^{i},\ldots,\boldsymbol{x}_{j}^{i}\right\}$ (3)
where $j$ is the length of the time-series for $1\leq j<J^{i}$ where $J$ is the maximum time step with timestamps of measurements $\left[t_{1}^{i},t_{2}^{i},\ldots,t_{j}^{i}\right]$. $\boldsymbol{x}_{j}^{i}$ contains $M$ set of features $\left[x_{j,1}^{i},x_{j,2}^{i},\ldots,x_{j,M}^{i}\right]^{\top}$ for timestamp $t_{j}^{i}$.
We undertake estimation of the underlying cumulative risk by mainly optimizing the negative logarithmic likelihood of the joint distribution of the event time and outcome with right-censoring. For those patients who have suffered the event, we capture both the outcome and the time at which it occurs. For censored patients, we capture the censoring time conditioned on the measurements recorded prior to the censoring. If we assume $\hat{a}_{t}=P\left(T=t\mid\mathcal{X}\right)$ (the output of the last layer node of the neural network module) represents the estimated probability of experiencing an event at time $t$, then the loss can be represented as
@@
where $\mathbb{1}$ is the indicator function. The first term represents optimising for uncensored patients and the second for those censored at the last noted time ie. alive. The estimated cumulative incidence risk, $\hat{F}$, is the probability that the patient dies at time $t$ or before, conditioned on all previous longitudinal measurements. By optimising for this loss, we estimate the actual risk distribution for each patient and a prediction can be made for arbitrary times across event times as outputs of the last layer.
As data are complex (i.e., static and time-varying), we seek to learn a latent representation using a conditional variational autoencoder that would improve learning for the task of survival analysis. There are two loss terms, namely the reconstruction (mean squared error) and the Kullback–Leibler (KL) divergence in the variational autoencoder. Estimating the distribution of the underlying latent factors relies on minimising the KL divergence between an approximation of the true posterior and the true distribution both of which are assumed to be multivariate Gaussians. As such, learning these distributions means optimising for their parameters, the mean and standard deviation. $z$ is the sampled latent vector from the probabilistic encoder for the learned Gaussian distribution with mean $\mu$ and standard deviation $\sigma$. For training, however, since sampling is a stochastic process, we use the reparameterization trick to backpropagate the gradient and represent the latent vector as the sum of a deterministic variable and an auxiliary independent random variable $\varepsilon$ [26]
@@
Figure 5 in the Appendix shows the structure of the proposed model. DySurv leverages the dynamic nature of deep learning time-series models combined with conditional variational autoencoders for multi-task learning of survival analysis risk prediction extending beyond the classical fixed binary event prediction from traditional machine learning models. Figure 1 shows the framework employed here built around DySurv to support robust risk estimation for clinical events such as death. Using a simple autoencoder has been shown to lead to overfitting and imbalanced learning of the reconstruction task that could harm learning the survival task whereas a VAE’s objective function is based on the reconstruction loss from a randomly sampled vector allowing for more robustness [12]. We concatenate the feature vectors from the static and time-series features together before feeding them into an encoder equipped with a Long-Short Term Memory (LSTM) cell.
The time-series components are compressed into a latent representation that is then used, as is, as input to the estimation of the negative log-likelihood loss function. A detailed description of the architecture can be found in the Appendix. Once the latent vector is sampled from the Gaussian distribution defined by these parameters, the lower-dimensional latent factors are used as input for an MLP module to optimize the survival task. The hyperparameters optimised in the network through grid search using the training and validation set included learning rate, batch size, $\alpha$, and dropout proportion. To minimise overfitting, we employ early stopping techniques.
@@
We compare the performance of our model to a collection of survival analysis models summarised in Table I and described in detail in the Appendix.

### v3

@@ Abstract
Objective: Machine learning applications for longitudinal electronic health records often forecast the risk of events at fixed time points whereas survival analysis achieves dynamic risk prediction by estimating time-to-event distributions. Here, we propose a novel conditional variational autoencoder-based method, DySurv, which uses a combination of static and longitudinal measurements from electronic health records to estimate the individual risk of death dynamically.
Materials and Methods: DySurv directly estimates the cumulative risk incidence function without making any parametric assumptions on the underlying stochastic process of the time-to-event. We evaluate DySurv on 6 time-to-event benchmark datasets in healthcare, as well as two real-world ICU EHR datasets extracted from eICU and MIMIC-IV.
Results: DySurv outperforms other existing statistical and deep learning approaches to time-to-event analysis across concordance and other metrics. It achieves time-dependent concordance of over 60% in the eICU case. It is also over 12% more accurate and 22% more sensitive than in-use ICU scores like APACHE and SOFA. The predictive capacity of DySurv is consistent and the survival estimates remain disentangled across different datasets.
Discussion: Our interdisciplinary framework successfully incorporates deep learning, survival analysis, and intensive care to create a novel method for time-to-event prediction from longitudinal health records. We test our method on several held-out test sets from a variety of healthcare datasets and compare it to existing in-use clinical risk scoring benchmarks.
Conclusion: While our method leverages non-parametric extensions to deep learning-guided estimations of the survival distribution, further deep learning paradigms could be explored.
@@ I BACKGROUND
Survival analysis refers to statistical approaches to estimating distributions of event times or times it takes for an event to happen as well as rates of survival over time while accounting for censoring. The events in question can be machine failures in industry or the occurrence of specific diseases and death [1]. In clinical practice, survival analysis can play a key role and provide valuable information to predict patient outcomes and guide treatment decisions [2]. While most traditional applications of survival analysis occur in epidemiology at the population level, with the rise of deep learning techniques, personalised estimation of survival times for individual patients has become possible [3]. However, the limitations of standard statistical models such as the Cox proportional hazards model include lack of complexity, constraining assumptions about the behaviour or proportions of covariate effects over time, and the reliance on only using static covariates [4]. While there are extensions of Cox to time-varying covariates, these still suffer from the proportionality assumption (often violated in practice) and are outperformed by simple deep learning models [5]. The proportional hazards assumption inherent in Cox models states that the hazard rates of different patients over time remain in fixed proportion to each other, in other words, despite the potential change in risk trajectories for a patient in the future, these changes remain fixed relative to other patients. Thus, implying that the hazard in measuring the effect of any predictor is constant over time. This severely restraints the predicted risk trajectories and they often do not capture real changes in risk over time for a patient [6]. In settings like healthcare, high-dimensional longitudinal information on the patient’s state can be an informative source for the prediction of clinical risk of mortality and other events.
@@
To incorporate longitudinal, high-dimensional, and potentially multimodal data and to move away from constraining assumptions of simple statistical models, deep learning methods have been proposed in survival analysis. These include extensions to the Cox model which estimate the parameters using a deep learning model. However, these models remain restrained with the proportionality assumption as Cox, such as the case of DeepSurv [7]. Existing methods often rely on parametric or semi-parametric assumptions of the survival distribution, thereby potentially restraining the predicted distribution away from a closer representation of the true survival distribution [8]. Other deep learning models include DeepHit and its extension to longitudinal data, Dynamic-DeepHit, which implements a custom loss and a simple recurrent neural network to avoid making any parametric assumptions on the survival distribution [9, 10]. Simple extensions of recurrent neural networks for estimating survival distributions have also been proposed [11, 12]. More recently, the autoencoder structure has been proposed to learn from a combination of data modalities by finding a lower-dimensional latent representation of the data. ConcatAE uses a simple autoencoder on multi-omics data to find a hidden representation then used in a multi-class classification task to predict discrete survival times [13]. This treats the survival analysis task as one of simple classification and does not aim to estimate the survival distribution at all. On the other hand, the variational autoencoder present in VAECox, directly estimates the survival distribution using the Cox loss function while taking advantage of the variational inference over traditional autoencoders [14]. Due to the reliance on the Cox loss, VAECox suffers from the often-violated proportionality assumption meaning its results in real-world datasets could be limited.
@@
Our work addresses these limitations by using a cumulative incidence risk estimation loss function based on the negative log-likelihood which requires making no parametric or proportionality assumptions on the survival distribution and hazard risks. We also extend the variational setting to conditional VAE which we find improves predictive performance. Our model, DySurv, is compatible with both static and longitudinal time-series data enabling more comprehensive learning from patient electronic health records. We validate our approach both in static and time-varying settings using six benchmark datasets as well as two large public open-access ICU datasets.The critical care setting provides high-frequency longitudinal measurements and short-term health outcomes in which dynamic risk stratification can help with urgent prognostication and prevention over just detection [15]. We aim to show that conditional variational inference and autoencoder reconstruction tasks can improve learning from complex time-series data by extracting latent features to optimise the survival task. DySurv accomplishes this by adding a conditional variational inference loss to the cumulative risk loss estimation from logistic hazards which aids in learning the survival task. Since DySurv is not only multimodal in input but also in output with both reconstruction and survival tasks included, conditional variational autoencoders (CVAE) provide better predictive performance [16]. By mapping high-dimensional covariates into a lower-dimensional space, CVAEs can effectively capture the underlying latent structure of the data to estimate survival outcomes. This capability allows deeper insights into dynamic mortality risk prediction in healthcare.
II METHODS
@@ II-A Data
Standard benchmark datasets contain only static features but here we implement survival analysis on the ICU datasets from MIMIC-IV and eICU which contain both static and time-series data. To show performance across different datasets and with different sizes, we will succinctly introduce these datasets. Across all datasets, the event in question is death. The datasets were split into 60% training, and 20% each for validation and testing. Quantile transformations have been applied for standardisation and fit only on the training dataset. Please see the Supplementary for a full list of features for each dataset.
@@ II-A1 SUPPORT
The Study to Understand Prognoses and Preferences for Outcomes and Risks of Treatments contains data from five care academic centres in the United States for patient survival in the following six months [17]. The result of the study was a prognostic model to estimate survival for seriously ill hospitalised patients. The dataset consists of 8,873 samples and 14 static features. A total of 68.1 percent of patients died with a median death time of 58 days.
@@ II-A2 METABRIC
The Molecular Taxonomy of Breast Cancer International Consortium contains genetic and clinical data from breast cancer patients with 1,904 samples and 9 static features [18]. 57.7% have an observed death with a median survival time of 116 months.
@@ II-A3 GBSG
The Rotterdam & German Breast Cancer Study Group contains treatment and clinical data on 2,232 breast cancer patients with 6 static features [19]. 79.6% of patients have an observed event.
@@ II-A4 NWTCO
The National Wilm’s Tumor dataset contains staging and clinical data on 4,028 Wilms’ tumour patients with 6 static features [20]. 16.0% of patients have an observed event.
@@
The simulated dataset contains discrete event times with 44 static features and 100,000 samples [21]. 37% of samples have been censored.
@@
The simulated dataset contains discrete event times with 5 static features and 50,000 samples [22]. 37% of samples have been censored.
@@ II-A7 MIMIC IV and eICU
We conduct experiments on the de-identified real-world ICU dataset Medical Information Mart for Intensive Care (MIMIC-IV v. 2.0, July 2022) including discharge information for more than 15,000 additional ICU patients compared to the previous release [23]. The dataset contains data from Beth Israel Deaconess Medical Center collected between 2008 and 2019. The dataset contains 71,935 samples of ICU stays with 33 static features (categorical features were one-hot encoded) and 65 time-varying features. Furthermore, we tested survival analysis models in the eICU Collaboration Research Database [24]. The eICU database was processed using postgreSQL and the pandas package. eICU is a multi-centre ICU database with over 200,859 patient unit encounters for 139,367 unique patients admitted between 2014 and 2015 to one of 335 ICUs at 208 hospitals located throughout the United States. The database is de-identified and includes vital sign measurements, demographic data, and diagnosis information. Static variables include age, sex, admission unit, and others that did not have missingness. For the time-series variables, we use forward filling as clinicians in practice would only consider the last recorded measurement and as has been done by previous work on these datasets [25, 26]. If the first set of measurements is missing for some time-varying features, instead of dropping those features or patients, we backward fill from the closest measurement in the future. 10.0% of patients had an observed event of death. The pre-processing of MIMIC-IV and eICU follows from previously published work on these datasets but is adapted for the survival scenario with the duration of stay in the ICU or the maximum time horizon for the event times defined as 10 days and time-series features taken in 72-hour timesteps after resampling to 1-hour intervals [27]. 9.5% of patients had an observed event of death in this cohort. Censoring, as always, is defined as loss-to-follow-up or discharge.
@@ II-B DySurv
${f(t)=-\frac{dF}{dt}}$ (2)
and $F(t)$ corresponds to the cumulative risk or incidence function $F(t)=\operatorname{P}(T\leq t)$. $T$ represents the time of the event, $\operatorname{P}$ is the probability, and $t$ is the specitic timestamp for risk estimation. The key to training a deep learning model is to learn an estimate of the cumulative incidence or risk function $\hat{F}(t)$ as the joint distribution of the event time and outcome label given the observations. In general, the time-to-event values can be left to be continuous depending on the model being considered but we discretise the time set into 10 equally spaced time periods for each dataset in the fashion of DeepHit and Dynamic-DeepHit [10]. As we discretise the time into intervals, we can estimate this event probability across arbitrary periods and remain faithful to the original survival analysis formulation rather than resorting to chained binary classification [28]. Furthermore, measurements are often right-censored, meaning that patients can leave the study or be lost to follow-up and they therefore do not experience an event at all. A common assumption is that this censoring is not important and independent of the outcome of the study itself [29]. Once we have an estimate of the risk, we can then simply obtain the survival function and curves by subtracting the cumulative risk from 1 as shown in equation 1.
@@
with $\mathcal{X}^{i}$ representing the feature matrix of patient $i$, $\mathcal{T}$ being the time-to-event i.e., the minimum of the event and censoring time, $y\in\{\varnothing,1\}$ being the label for the outcome with $\varnothing$ representing right-censoring, and $N$ samples included. Static features have been expanded into time-series by replication padding matching the respective time-series component. Effectively, the feature matrix then consists only of time-series features. A feature matrix for patient $i$, $\mathcal{X}^{i}$, can be seen as
$\mathcal{X}^{i}=\left\{\bm{x}_{1}^{i},\bm{x}_{2}^{i},\ldots,\bm{x}_{j}^{i}\right\}$ (4)
where $j$ is the length of the time-series for $1\leq j<J^{i}$ where $J$ is the maximum time step with timestamps of measurements $\left[t_{1}^{i},t_{2}^{i},\ldots,t_{j}^{i}\right]$. $\bm{x}_{j}^{i}$ contains $M$ set of features $\left[x_{j,1}^{i},x_{j,2}^{i},\ldots,x_{j,M}^{i}\right]^{\top}$ for timestamp $t_{j}^{i}$.
We undertake estimation of the underlying cumulative risk by mainly optimizing the negative logarithmic likelihood of the joint distribution of the event time and outcome with right-censoring. For those patients who have suffered the event, we capture both the outcome and the time at which it occurs. For censored patients, we capture the censoring time conditioned on the measurements recorded prior to the censoring. If we assume $\hat{a}_{t}=P\left(T=t\mid\mathcal{X}\right)$ represents the estimated probability of experiencing an event at time $t$, then the loss can be represented as
@@
where $\mathbb{1}$ is the indicator function. The first term represents optimising for cumulative incidence risk accounting for uncensored patients and the second for those censored at the last noted time ie. alive. The estimated cumulative incidence risk, $\hat{F}$, is the probability that the patient dies at time $t$ or before, conditioned on all previous longitudinal measurements. By optimising for this loss, we estimate the actual risk distribution for each patient and a prediction can be made for arbitrary times across event times.
As data are complex (i.e., static and time-varying), we seek to learn a latent representation using a conditional variational autoencoder that would improve learning for the task of survival analysis [31] [^1] For those readers less familiar with the autoencoder architecture and machine learning background for survival analysis, we recommend the following insightful surveys and chapters to peruse: [8, 32, 33]. Since we are working with a deep learning model, our goal is to minimise a specific loss function relevant to our task. The loss will penalise our model when it starts learning away from the task in question. The first loss term relates to learning an efficient cumulative risk distribution to represent survival as described above in Equation 5. There are two further loss terms, namely the reconstruction (mean squared error) and the Kullback–Leibler (KL) divergence in the variational autoencoder [34]. The first helps the model learn better representations of the input data, while the second helps learn better latent representations. Estimating the distribution of the underlying latent factors relies on minimising the KL divergence between an approximation of the true posterior and the true distribution both of which are assumed to be multivariate Gaussians. As such, learning these distributions means optimising for their parameters, the mean and standard deviation. $z$ is the sampled latent vector from the probabilistic encoder for the learned Gaussian distribution with mean $\mu$ and standard deviation $\sigma$. For training, however, since sampling is a stochastic process, we use the reparameterization trick to backpropagate the gradient and represent the latent vector as the sum of a deterministic variable and an auxiliary independent random variable $\varepsilon$ [35]
@@
DySurv leverages the dynamic nature of deep learning time-series models combined with conditional variational autoencoders for multi-task learning of survival analysis risk prediction extending beyond the classical fixed binary event prediction from traditional machine learning models. Figure 1 shows the framework employed here built around DySurv to support robust risk estimation for clinical events such as death. Using a simple autoencoder has been shown to lead to overfitting and imbalanced learning of the reconstruction task that could harm learning the survival task whereas a VAE’s objective function is based on the reconstruction loss from a randomly sampled vector allowing for more robustness [14]. We concatenate the feature vectors from the static and time-series features together before feeding them into an encoder equipped with a Long-short-term-memory (LSTM) cell. The LSTM module allows us to learn the temporal patterns in the longitudinal data and represent it in a hidden embedding for downstream tasks such as the estimation of the survival function. A short primer on the LSTM cell can be found in Supplementary Section C.
The time-series components are compressed into a latent representation that is then used, as is, as input to the estimation of the negative log-likelihood loss function. A detailed description of the architecture can be found in the Supplementary. Once the latent vector is sampled from the Gaussian distribution defined by these parameters, the lower-dimensional latent factors are used as input for an MLP module to optimize the survival task. In the case of the ICU datasets, for example, each output node of this MLP, a representation of $\hat{a}_{t}$, then corresponds to a 1-day risk prediction in the ICU for a maximum time horizon of 10 days. The hyperparameters optimised in the network through grid search using the training and validation set included learning rate, batch size, $\alpha$, and dropout proportion. To minimise overfitting, we employ early stopping.
@@
We compare the performance of our model to a collection of survival analysis models summarised in Table I and described in detail in the Supplementary. Unlike DySurv, existing methods like Cox-based methods as well as some deep learning alternatives, rely on assuming a specific statistical distribution of the survival distribution, thus making parametric distributions that restrict their predictive power. This assumption is described in further detail in Supplementary Section B.